### v1

# Kernel Ridge Regression for Efficient Learning of High-Capacity Hopfield Networks

[…]
Kernel Ridge Regression for Efficient Learning of High-Capacity Hopfield Networks
(2015; 2015)
keywords:
Hopfield network, Kernel ridge regression, Associative memory, Storage capacity, Closed-form solution, Efficient learning.
{summary}
Hebbian learning limits Hopfield network capacity. While kernel methods like Kernel Logistic Regression (KLR) improve performance via iterative learning, we propose Kernel Ridge Regression (KRR) as an alternative. KRR learns dual variables non-iteratively via a closed-form solution, offering significant learning speed advantages. We show KRR achieves comparably high storage capacity (reaching ratio 1.5 shown) and noise robustness (recalling from around 80% corrupted patterns) as KLR, while drastically reducing training time, establishing KRR as an efficient method for building high-performance associative memories.
1 Introduction
Hopfield networks [1] model content-addressable memory but suffer from limited storage capacity ($\beta=P/N\approx 0.14$, where $P$ is the number of patterns and $N$ is the number of neurons) when using the standard Hebbian rule [2], leading to retrieval errors for higher loads. Various approaches have sought to improve capacity, often framing learning as a supervised task to ensure patterns are stable fixed points. While linear methods like LLR [3] offer moderate improvements, they are constrained by linear separability. Non-linear kernel methods show greater promise; Kernel Logistic Regression (KLR) recently demonstrated dramatically enhanced capacity and robustness via iterative learning. However, this iterative process can be computationally intensive [4].
In this letter, we propose Kernel Ridge Regression (KRR) [5] as an alternative, efficient kernel-based approach for Hopfield networks. KRR utilizes the kernel trick but predicts bipolar states via regression with a squared-error loss, crucially offering a non-iterative, closed-form learning solution. We implement and evaluate KRR, comparing its capacity and robustness against Hebbian, LLR, and KLR. Our aim is to verify if KRR can match KLR’s high performance while offering substantial training speed advantages. Our results confirm KRR achieves comparable state-of-the-art performance with significantly reduced learning time, establishing it as a potent and efficient training method.
This work provides the first (or among the first) empirical demonstration comparing KRR with KLR in the context of Hopfield networks, highlighting KRR’s significant efficiency gains.
2 Methods
2.1 Model Setup
We consider a network of $N$ bipolar neurons $\{-1,1\}^{N}$. Let $\{\boldsymbol{\xi}^{\mu}\}_{\mu=1}^{P}$ be the set of $P$ bipolar patterns to be stored in the network’s associative memory. For supervised learning approaches requiring binary targets (LLR and KLR), we transform these patterns into target vectors $\mathbf{t}^{\mu}\in\{0,1\}^{N}$ where $t_{i}^{\mu}=(\xi_{i}^{\mu}+1)/2$. For KRR, the original bipolar values $\xi_{i}^{\mu}$ are used as targets. The network state at discrete time $t$ is represented by the vector $\mathbf{s}(t)\in\{-1,1\}^{N}$. This state evolves over subsequent time steps according to the dynamics determined by the specific learning algorithm employed (detailed in Sec. 2.2 and 2.3).
2.2 Learning Algorithms
We compare the proposed KRR method against several baseline and related learning approaches.
2.2.1 Baselines
Hebbian Learning:
The standard Hebbian rule [1, 2] computes the weight matrix $\mathbf{W}^{\text{Heb}}=(1/N)\mathbf{X}^{\top}\mathbf{X}$ (diagonals zeroed). Recall uses $\mathrm{s}_{i}(t+1)=\text{sign}(\sum_{j\neq i}\mathbf{W}_{ij}^{\text{Heb}}\mathbf{s}_{j}(t))$ This serves as the classical baseline.
Linear Logistic Regression (LLR):
A linear supervised baseline [3], where weights $\mathbf{W}^{\text{LLR}}$ are learned iteratively via gradient descent to optimize logistic loss with a linear predictor for predicting neuron states. Recall uses the same update rule as Hebbian, with $\mathbf{W}^{\text{LLR}}$.
Kernel Logistic Regression (KLR):
An iterative kernel-based method shown to achieve high capacity [4]. It learns dual variables $\boldsymbol{\alpha}^{\text{KLR}}$ by iteratively optimizing regularized logistic loss. Recall involves kernel evaluations (Sec. 2.3). We use results from [4] for comparison.
2.2.2 Kernel Ridge Regression Learning
As the primary focus of this work, we propose Kernel Ridge Regression (KRR) [5] as an efficient supervised learning approach. KRR frames the task as regression, aiming to directly predict the target bipolar state $\xi_{i}^{\mu}\in\{0,1\}$ for each neuron $i$.
Leveraging the kernel trick with a chosen kernel function $K(\cdot,\cdot)$, KRR finds a predictor function of the form $f_{i}(\boldsymbol{\xi}^{\mu})=\sum_{\mu=1}^{P}K(\boldsymbol{\xi}^{\nu},\boldsymbol{\xi}^{\mu})\alpha_{\mu i}$. This function minimizes the sum of squared errors between the predictions $f_{i}(\boldsymbol{\xi}^{\mu})$ and the targets $\xi_{i}^{\nu}$ over all patterns $\nu$, subject to L2 regularization on the function norm in the reproducing kernel Hilbert space.
A key advantage of KRR is that the optimal dual variables $\boldsymbol{\alpha}_{i}=(\alpha_{1i},\ldots,\alpha_{Pi})^{\top}$ for neuron $i$ can be obtained non-iteratively via the following closed-form solution:
$\boldsymbol{\alpha}_{i}=(\mathbf{K}+\lambda\mathbf{I})^{-1}\mathbf{y}_{i}$ (1)
where $\mathbf{K}$ is the $P\times P$ kernel matrix ($\mathbf{K}_{\nu\mu}=K(\boldsymbol{\xi}^{\nu},\boldsymbol{\xi}^{\mu})$), $\mathbf{y}_{i}=(\xi_{i}^{1},\ldots,\xi_{i}^{P})^{\top}$ is the target vector containing the bipolar states for neuron $i$ across all patterns, $\mathbf{I}$ is the $P\times P$ identity matrix, $\lambda>0$ is the regularization parameter.
The entire $P\times N$ dual variable matrix $\boldsymbol{\alpha}$ can thus be computed efficiently by solving the linear system $(\mathbf{K}+\lambda\mathbf{I})\boldsymbol{\alpha}=\mathbf{X}$, where $\mathbf{X}$ is the $P\times N$ matrix of stored patterns. This eliminates the need for iterative optimization required by LLR and KLR.
2.3 Recall Process
The state update rule differs between the baseline/linear methods and the kernel methods. After learning, the network state $\mathbf{s}(t)\in\{-1,1\}^{N}$ evolves over discrete time steps $T$.
Hebbian and LLR:
$s_{i}(t)=\text{sign}(\sum_{j\neq i}\mathbf{W}_{ij}s_{j}(t))$ (using $\mathbf{W}^{\text{Heb}}$ or $\mathbf{W}^{\text{Heb}}$), where $s_{i}(t)$ is the $i$-th element of the state $\mathbf{s}(t)$.
KLR and KRR:
Recall operates without an explicit $N\times N$ weight matrix $\mathbf{W}$, instead utilizing the learned $P\times N$ dual variable matrix $\boldsymbol{\alpha}$ and the stored patterns $\mathbf{X}$. The update rule involves three steps, performed synchronously for all neurons:
  1. 1.
    Compute kernel values:
    $\mathbf{k}_{\mathbf{s}(t)}=\left[K(\mathbf{s}(t),\boldsymbol{\xi}^{1}),\ldots,K(\mathbf{s}(t),\boldsymbol{\xi}^{P})\right]$ (size $1\times P$).
  2. 2.
    Compute activation potential: $\mathbf{h}(\mathbf{s(t)})=\mathbf{k}_{\mathbf{s}(t)}\boldsymbol{\alpha}$ (size $1\times N$).
  3. 3.
    Update the state: Determine the next state using the sign function. We set the activation threshold $\boldsymbol{\theta}$ to zero for simplicity: $\mathbf{s}(t+1)=\text{sign}(\mathbf{h}(\mathbf{s(t)}))$ (size $1\times N$).
It is important to note that KLR/KRR recall involves kernel computations between the current state and all $P$ stored patterns at each update step.
3 Experiments
3.1 Experimental Setup
We conducted simulations to evaluate the performance of the proposed KRR approach and compare it against baseline methods. All simulations utilized networks composed of $N=500$ bipolar neurons. Random bipolar patterns $\boldsymbol{\xi}^{\mu}\in\{0,1\}^{N}$ were generated for storage, with each element chosen independently and with equal probability $P(\xi_{i}^{\mu}=1)=0.5$.
We compared four learning algorithms:
Hebbian:
Standard Hebbian learning rule [1, 2].
LLR:
Linear Logistic Regression, implemented as described in [3], serving as a linear supervised baseline.
KLR:
Kernel Logistic Regression, using the implementation and results reported in our previous work [4] for direct comparison.
KRR:
The proposed Kernel Ridge Regression approach (Sec. 2.2.2).
For both KLR and KRR, we employed the Radial Basis Function (RBF) kernel $K(\mathbf{x},\mathbf{y})=\exp(\gamma\|\mathbf{x}-\mathbf{y}\|^{2})$ with the kernel width parameter set to $\gamma=1/N$. Based on preliminary experiments and findings in [4], the L2 regularization parameter was set to $\lambda=0.01$ for both LLR, KLR and KRR, as this value provided robust performance. For the iterative learning methods (LLR and KLR), the learning rate was $\eta=0.1$ and the number of learning updates was fixed at 200, ensuring sufficient convergence in our previous study [4].
3.2 Storage Capacity and Noise Robustness Evaluation
We first assessed the storage capacity of each learning algorithm by measuring the recall success rate as a function of the storage load $\beta=P/N$. Recall was initiated from the original stored patterns ($\mathbf{s}(0)=\boldsymbol{\xi}^{\mu}$), testing the network’s ability to maintain them as stable fixed points. Figure 1 plots the success rate against $\beta$ for $N=500$ neurons. As expected, the standard Hebbian network exhibits a sharp performance drop around the theoretical limit of $\beta\approx 0.14$ [2]. LLR significantly improves upon this, maintaining near-perfect recall up to $\beta\approx 0.85$, but fails completely as the load approaches $\beta=0.95$. In stark contrast, both KLR, consistent with our previous findings [4], and the proposed KRR demonstrate substantially higher capacity. Both kernel methods achieved a 100% success rate across the entire tested range, including loads exceeding $\beta=1.0$ (up to $\beta=1.5$ shown), indicating their capability to reliably store patterns even when $P>N$.
Next, we evaluated noise robustness by examining the size of the basins of attraction. We fixed the storage load at an intermediate level, $\beta=0.2$ ($P=100$ for $N=500$), where both LLR and kernel methods typically succeed from clean initial states, but Hebbian learning fails. We initiated recall from corrupted versions of a target pattern, varying the initial overlap $m(0)$. Figure 2 shows the mean final overlap $m(T)$ achieved after $T=25$ steps as a function of $m(0)$. The Hebbian network fails to converge to the target pattern regardless of the initial overlap within this range. LLR shows improved robustness, successfully recalling the pattern ($m(T)\approx 1.0$) if the initial overlap $m(0)$ is above approximately 0.4. Both KLR and KRR exhibit significantly enhanced robustness. They achieve perfect recall ($m(T)\approx 1.0$) even when starting from much noisier states, requiring an initial overlap $m(0)$ of only about 0.2. This demonstrates that both kernel methods create considerably larger basins of attraction compared to LLR and, especially, the Hebbian rule, making them more resilient to input noise. Notably, the robustness profiles of KLR and KRR appear virtually identical under these conditions.
Figure 1: Recall success rate vs. storage load ($\beta=P/N$) for Hebbian, LLR, and KRR ($N=500$).
Figure 2: Noise robustness comparison: Final Overlap $m(T)$ vs. Initial Overlap $m(0)$ at $\beta=0.2$ for Hebbian, LLR, and KRR.
3.3 Learning Time Comparison
To quantitatively assess the practical efficiency trade-offs alongside the performance benefits, we measured and compared the learning phase computational time for LLR, KLR, and KRR across various storage loads ($\beta$). Wall-clock time was measured using Python’s time.time() function, capturing the time required to complete the learning process for each method on a standard desktop CPU. The results, averaged over 3 trials for $N=500$, are presented in Table 1 and visualized in Figure 3 (note the logarithmic scale for the time axis).
Hebbian learning, being non-iterative and computationally simple ($O(N^{2})$), completes extremely quickly. Its learning time is omitted from the comparison table/figure for clarity, as it is significantly shorter than the other methods. The timing results reveal significant differences in efficiency and scaling among the supervised methods. LLR, requiring 100 iterative updates in our setup, exhibited the slowest performance, with learning time increasing approximately linearly with the storage load $\beta$ (from around 3 seconds at $\beta=0.1$ to around 30 seconds at $\beta=1.0$. This scaling is consistent with its need to process $P$ patterns across $N$ neurons iteratively.
KLR, using 200 iterative updates, was considerably faster than LLR across all tested loads. Its learning time, however, still showed a clear dependence on $\beta$, increasing from around 0.07 seconds at $\beta=0.1$ to around 1.35 seconds at $\beta=1.0$. This increase is likely dominated by computations involving the $P\times P$ kernel matrix within the iterative optimization process.
Crucially, KRR demonstrated outstanding learning efficiency. Leveraging its non-iterative, closed-form solution (Eq. 1), KRR consistently achieved the fastest learning times by a large margin. Its learning time remained under 0.1 seconds even at the maximum load of $\beta=1.0$ ( 0.07 seconds), showing only a very modest increase with $\beta$. This time is primarily composed of the one-time kernel matrix construction and the subsequent solving of the linear system. At $\beta=1.0$, KRR was approximately 20 times faster than KLR and over 400 times faster than LLR, highlighting the dramatic speed advantage conferred by its closed-form solution.
These results underscore the significant practical benefit of KRR for training Hopfield networks. While both KLR and KRR deliver comparable, state-of-the-art storage performance, KRR offers substantially higher training efficiency, especially at higher storage loads. This makes KRR a highly attractive option, particularly when computational resources for training are limited or when models need to be trained rapidly. The scalability challenge for recall ($O(PN)$ complexity), however, remains similar for both KLR and KRR.
Table 1: Mean learning time (seconds) for different algorithms and storage loads ($\beta=P/N$) with $N=500$. Values shown are mean $\pm$ standard deviation over 3 trials.
Storage Load Learning Algorithm
$\beta$ LLR (100 Updates) KLR (200 Updates) KRR (Closed-form)
0.1 $3.010\pm 0.045$ $0.072\pm 0.014$ $0.009\pm 0.001$
0.2 $5.988\pm 0.128$ $0.188\pm 0.031$ $0.010\pm 0.003$
0.4 $12.180\pm 0.090$ $0.365\pm 0.009$ $0.026\pm 0.003$
0.6 $18.356\pm 0.426$ $0.723\pm 0.042$ $0.035\pm 0.005$
0.8 $24.350\pm 0.310$ $0.804\pm 0.010$ $0.066\pm 0.017$
1.0 $30.363\pm 0.844$ $1.350\pm 0.088$ $0.069\pm 0.003$
Figure 3: Learning time comparison (log scale) vs. storage load ($\beta$) for LLR, KLR, and KRR ($N=500$, avg. over 3 trials).
4 Discussion
Our comprehensive experiments, conducted with $N=500$ neurons, unequivocally demonstrate that applying kernel-based supervised learning dramatically enhances the performance of Hopfield networks compared to the classical Hebbian rule and linear supervised methods like Linear Logistic Regression (LLR). Both Kernel Logistic Regression (KLR) [4] and the proposed Kernel Ridge Regression (KRR) exhibited substantial improvements in storage capacity and noise robustness.
The gains in storage capacity are particularly remarkable. Both KLR and KRR achieved perfect recall success rates up to $\beta=1.5$ (Fig. 1), vastly exceeding the Hebbian limit ($\beta\approx 0.14$[2] and also significantly surpassing LLR (which failed around $\beta\approx 0.9$). The ability of these kernel methods to operate flawlessly even when $P>N$ highlights their powerful representational capability, likely stemming from the effective pattern separation enabled by the implicit high-dimensional mapping via the RBF kernel.
In terms of noise robustness, KLR and KRR also showed marked advantages (Fig. 2). They formed considerably larger basins of attraction, successfully recalling patterns from initial states with much lower overlap ($m(0)\approx 0.2$) compared to LLR ($m(0)\approx 0.4$) and the Hebbian rule. This superior robustness underscores their effectiveness with corrupted inputs. Notably, the recall performance profiles of KLR and KRR were virtually identical in both tests, suggesting the choice between logistic and squared-error loss might not critically affect the final attractor landscape in this bipolar setting, using our chosen hyperparameters ($\gamma=1/N$, $\lambda=0.01$).
The most significant practical difference between KRR and KLR emerged in learning efficiency. KRR’s non-iterative, closed-form solution (Eq. 1) provided a dramatic speed advantage across all storage loads (Table 1, Fig. 3). At $\beta=1.0$, KRR learned approximately 20 times faster than KLR (200 updates) and over 400 times faster than LLR (100 updates). While KLR was also faster than LLR in our tests, KRR offers unparalleled training speed among these high-performing methods.
Despite KRR’s learning efficiency, the recall computational cost remains a crucial consideration, common to both KLR and KRR. The recall process requires $P$ kernel evaluations per step (Sec 2.3), scaling as $O(PN)$, unlike the $O(N^{2})$ complexity of Hebbian/LLR recall. This presents a scalability challenge for large $P$, especially in the high-capacity $P>N$ regime. This trade-off between high performance and recall cost motivates exploring kernel approximations like the Nyströem method [6].
This work aligns with the Kernel Memory Network framework [7], empirically validating KLR/KRR for achieving high capacity (even $P>N$) and highlighting KRR’s efficiency. Future directions include validating approximations like Nyströem, evaluating other kernel functions, developing online learning variants, and perhaps analyzing the attractor landscape more deeply.
5 Conclusion
We have demonstrated that Kernel Ridge Regression (KRR) provides a powerful and highly efficient learning mechanism for Hopfield-type associative memory networks. By leveraging kernel methods to implicitly perform non-linear feature mapping, KRR achieves exceptional storage capacity (successfully operating even for $P>N$) and noise robustness, comparable to state-of-the-art iterative methods like KLR. Critically, KRR achieves this high performance with significantly reduced learning time due to its non-iterative, closed-form solution, offering speedups of over an order of magnitude compared to KLR and several orders of magnitude compared to LLR in our tests. While the recall complexity scaling with $P$ remains a consideration for both KRR and KLR, KRR’s combination of top-tier storage performance and superior training efficiency makes it a particularly compelling and practical approach for building high-performance associative memory systems, especially when rapid model development or retraining is required. This work further highlights the potential of applying modern kernel-based machine learning techniques to enhance and understand classic neural network models.